OpenStudy (anonymous):

Evaluate S5 for 300 + 150 + 75 + …

OpenStudy (mizuki):

wat is the question??

OpenStudy (anonymous):

that is the question thats all it says and it gives me 4 answer choices

OpenStudy (amoodarya):

this is (sum of)geometric sequenc \[a+\frac{a}{2}+\frac{a}{4}+...=\frac{a}{1-\frac{1}{2}}\]

OpenStudy (amoodarya):